OpenStudy (anonymous):

What is the restriction on the quotient of quantity 5 x squared minus 10 x minus 15 divided by quantity x squared minus 9 divided by quantity 6 x plus 12 divided by quantity x plus 2?

OpenStudy (anonymous):

so ur equation is (5x^2 -10x -15)/(x^2 -9)/(6x+12)/(x^+2)?

OpenStudy (anonymous):

Yep, that's it:)

OpenStudy (anonymous):

How would I go about solving this?

OpenStudy (anonymous):

just by looking at the equation u know the top part can be 0 and all real numbers known as IR. The others can have possible restrictions if SET EQUAL TO ZERO because it is undefined when u a number for example 1/0.

OpenStudy (anonymous):

I'm a bit lost, how do I set these to zero?

OpenStudy (anonymous):

so really you have (5x^2-10x-15)/(x^2 -9)(6x+12)(x+2)

OpenStudy (anonymous):

all u care about is the restriction right. so set those () each = 0 and solve for x

OpenStudy (anonymous):

I get 'how' I would set these to zero, but how would I solve the parts with that? Aren't there no like terms.?

OpenStudy (campbell_st):

well you are dividing by a fraction so flip and multiply which gives \[\frac{5x^2 -10x -15}{x^2 -9} \times \frac{6x + 12}{x + 2}\] so to find the restrictions... I'd first factor the quadratics then write everything over a common denominator \[\frac{5(x +1)(x-3)6(x + 2)}{(x +3)(x-3)(x+2)}\] so the restrictions occur in the denominator... identify the values of x that make the denominator zero... there will be 2 points of discontinuity and 1 vertical asymptote

OpenStudy (campbell_st):

depending on the way it should be answered the expression above can be simplified to \[\frac{30(x + 1)}{(x +3)}\]

OpenStudy (campbell_st):

by stating the restrictions on the simplified form you will only get 1 vertical asymptote...

OpenStudy (anonymous):

ok, so I then cancel some of the terms. I was confused by this because they want me to find the restriction of the product.? Would it be -3 since that's the opposite of x + 3?

OpenStudy (campbell_st):

well there are actually 3 restrictions.... 2 are points of discontinuity... there occur in the original product... but not in the simplified form. there is 1 restriction that occurs in the original and simplified form... the question doesn't state how many answers ... but for me... I'd provide all 3 ...

OpenStudy (anonymous):

-3 would make the the equation be 0/0 which is ok

OpenStudy (campbell_st):

x = -3 doesn't but x = 3 does...

OpenStudy (campbell_st):

when you get 0/0 you get a point of discontinuity

OpenStudy (campbell_st):

here is a nice summary http://www.math.brown.edu/UTRA/discontinuities.html

OpenStudy (anonymous):

so is it -3 or 3? I have both options..

OpenStudy (anonymous):

plug what u find into the top part of orginal equation because u know the bottom is ?/0

OpenStudy (campbell_st):

well there are both restictions... its just determining their type... there is 1 other restiction in the domain

OpenStudy (anonymous):

If the denominator was x + 3, I would think that it would be the opposite of -3, but I'm not that good. My options are: x ≠ -2 x ≠ -3 x ≠ 3 x ≠ 9

OpenStudy (campbell_st):

oops slight typo should be \[\frac{5(x +1)(x-3)(x +2)}{(x-3)(x+3)6(x+2)}\] so cancel the common factors... and look at the denominator restriction...

OpenStudy (anonymous):

x value = -3 plug into your equation check to see if it is undefined

OpenStudy (campbell_st):

so simplifying you get \[\frac{5(x +1)}{6(x +3)}\] the restriction occurs when x + 3 = 0 hope it helps

OpenStudy (anonymous):

Ok, well, thanks, both of you for your help:)

OpenStudy (anonymous):

yw just hope ur not more confused and u understand it better.

OpenStudy (anonymous):

I do, thx again, :) I just posted a similar one but is just asking for simplification if you have any time for one more.

OpenStudy (anonymous):

so 3 is the restriction?

OpenStudy (anonymous):

I think it would be negative three, because the restriction is generally the opposite. Not sure though.

OpenStudy (anonymous):

Yep, just finshed the test- the answer is negative 3